OpenStudy (anonymous):

pump a can empty a pool in 4 hours. pump b can empty the same pool in 12 hours. if both pumps are working together, which of the following is the amount of time it will take them to empty the pool

OpenStudy (oaktree):

 Now solve.

hero (hero):

\[ t = \frac{4 \times 12}{4 + 12}\]

OpenStudy (anonymous):

3 hours

hero (hero):

In general, for similar problems \[t = \frac{a \times b}{a + b}\] a = 1st Person b = 2nd Person

OpenStudy (anonymous):

thank you